OpenStudy (justanothersmith):

Simplify this expression 5a/11a? I don't know if it's in correct format but it's 5a over 11a. I don't know how to do this, I take online classes and I don't hardly get help.

OpenStudy (anonymous):

5/11 The a's divide out.

OpenStudy (justanothersmith):

Is it really as simple as that? o-o Well, I feel dumb now. Thank you.

ganeshie8 (ganeshie8):

It's not as simple as that if you think about why \(a\)'s must kill each other out

OpenStudy (justanothersmith):

They can't just kill eachother with swords and call it the results? (I'm sorry I attempt to create funny comments.)

ganeshie8 (ganeshie8):

Imagine you have 12 cookies, and you want to distribute them equally to your 4 friends

ganeshie8 (ganeshie8):

How many cookies each friend gets ?

OpenStudy (justanothersmith):

3?

ganeshie8 (ganeshie8):

How ?

OpenStudy (justanothersmith):

12 divided by 4 equals 3?

ganeshie8 (ganeshie8):

Exactly! So you can think of the "division operation" intuitively as distributing cookies to friends : \[\dfrac{12}{4}\] top is number of cookies you have bottom is number of friends you want to distribute the cookies to

ganeshie8 (ganeshie8):

The result "3" tells you the number of cookies EACH friend gets

OpenStudy (justanothersmith):

Unless I kill a friend and share their cookies right?

ganeshie8 (ganeshie8):

Unless you're partial to one of your friends.. it won't be an equal distribution anymore

ganeshie8 (ganeshie8):

Let me ask you a difficult question now

OpenStudy (anonymous):

(5a/a)/(11a/a) = 5/11

ganeshie8 (ganeshie8):

Imagine you have killed half of your friends and ate half of your cookies.

ganeshie8 (ganeshie8):

Then, how many cookies and friends you're left with ?

ganeshie8 (ganeshie8):

I don't like this idea of killing, maybe let me rephrase the question.. one moment

OpenStudy (justanothersmith):

Should we say they half of the friends are allergic to the cookies then?

ganeshie8 (ganeshie8):

Suppose you don't want to give cookies to half of your friends as you don't like them. So you ate half the cookies yourself and want to distribute the remaining half of cookies to the the remaining half of friends that you love so much.

ganeshie8 (ganeshie8):

I want you think about above situation. Now, after eating half of the cookies, if you distribute the remaining cookies equally to half of your friends, how many cookies does EACH of your friend get ?

OpenStudy (justanothersmith):

Lets see 12 cookies to start, 5 people(including myself) 2.4 cookies for everyone? o.o Okay, Let's take it to 2 cookies for everyone, I ate 2 cookies leaving 10, 2 friends are allergic and 2 friends ate their 2, that would would leave 6 cookies left and since only three could eat them that would give them three 2 more cookies meaning there would be 0 cookies, but 3 people got 4 cookies each. Right?

ganeshie8 (ganeshie8):

don't include yourself

ganeshie8 (ganeshie8):

your friends are the guests, you're just happy watching them eat

OpenStudy (justanothersmith):

Okay, so only 2 friends get to eat since the other 2 cannot? But in the questions you said I ate half of the cookies and that there were 12, so 2 people get 3 each and the other 2 don't get any due to allergies?

ganeshie8 (ganeshie8):

Right, so each of your loving friends are getting just 3 cookies again. recall that earlier also they got the same 3 cookies

ganeshie8 (ganeshie8):

Distributing 12 cookies to 4 friends is same as distributing 6 cookies to 2 friends

ganeshie8 (ganeshie8):

I mean, each friend gets the same number of cookies in both situations

ganeshie8 (ganeshie8):

Could you explain why they get the same number of cookies ?

OpenStudy (justanothersmith):

I mean, I don't get how 4 friends sharing 12 cookies and 2 friends sharing 6 cookies is even, but I am going to say because there is 12 cookies in both situations. I don't even know, It's 2:30AM and I have alot more "homework" to do...

ganeshie8 (ganeshie8):

It's okay, you're doing great and you know enough to complete your homework. Just think about this when you have lots of time : \[\dfrac{12}{4} = 3\] \[\dfrac{6}{2} = 3\] Why they both give the same answer ?

OpenStudy (justanothersmith):

Ah, I get that. I wasn't thinking of the part I ate half of the cookies leaving 6 between 2 friends. XD

ganeshie8 (ganeshie8):

More generally, think about why below "rule" is true: \[\dfrac{a}{b} = \dfrac{a/c}{b/c}\] The answer doesn't change if you divide top and bottom by the same number \(c\)

ganeshie8 (ganeshie8):

Awesome! I see you have nailed this division thingy! keep going with the rest of problems :)

OpenStudy (justanothersmith):

Well, I ran into another road block. It's asking me for the simplest form of 4(-m-1)-5. I am terrible at math as you can tell.

ganeshie8 (ganeshie8):

Could you take a screenshot of that expression and attach here if psble ?

OpenStudy (justanothersmith):

Sure. One second.

OpenStudy (justanothersmith):

It won't let me send once I attach a photo.

OpenStudy (justanothersmith):

This is it though. http://prntscr.com/b0gbc0

ganeshie8 (ganeshie8):

That's okay, so is it like this ? \[ \LARGE 4(-m-1)-5\]

ganeshie8 (ganeshie8):

Familiar with PEMDAS ?

OpenStudy (justanothersmith):

I am.

ganeshie8 (ganeshie8):

Good. So first look for any parenthesis in the expression and simplify the stuff inside them

ganeshie8 (ganeshie8):

\[ \LARGE 4(\color{red}{-m-1})-5\]

OpenStudy (justanothersmith):

So when you do that would inside the parenthesis equal -1m or 1m by a chance?

ganeshie8 (ganeshie8):

There is nothing to simplify inside the parenthesis

ganeshie8 (ganeshie8):

Get out of parenthesis and distribute 4

ganeshie8 (ganeshie8):

\[ \LARGE \color{red}{4(-m-1)}-5\]

OpenStudy (justanothersmith):

Meaning something like -4m -4 -5?

ganeshie8 (ganeshie8):

\[ \LARGE \color{red}{4(-m) + 4(-1)}-5\]

ganeshie8 (ganeshie8):

\[ \LARGE \color{red}{-4m - 4}-5\]

ganeshie8 (ganeshie8):

Yes, next combine the numbers -4 and -5

ganeshie8 (ganeshie8):

(-4) + (-5) = ?

OpenStudy (justanothersmith):

So -4m -9

ganeshie8 (ganeshie8):

Perfect!

OpenStudy (justanothersmith):

Awesome.

ganeshie8 (ganeshie8):

As you can see, these are simple if you understand few rules and follow them

OpenStudy (justanothersmith):

\[n(-4n^3-8)+7n^4\]

OpenStudy (justanothersmith):

So here I would distribute the n?

ganeshie8 (ganeshie8):

Yes

OpenStudy (justanothersmith):

\[-4n^4 -8n + 7n^4\]

ganeshie8 (ganeshie8):

Looks good, but it simplifies further

OpenStudy (justanothersmith):

I know I was taking it step by step. \[3n^4 - 8n\] Final?

ganeshie8 (ganeshie8):

Yep! can't mess with it further

OpenStudy (skullpatrol):

Always look for "similar" terms :-) That is two monomials that are exactly alike or are the same except for their numerical coefficients.

OpenStudy (justanothersmith):

Alright, Thank you all. :D

OpenStudy (skullpatrol):

Thanks for trying to learn and understand :D

ganeshie8 (ganeshie8):

Np :) You did a great job in understanding these tricky things ! keep up the good work !

OpenStudy (zzr0ck3r):

@ganeshie8 I just quoted you on my facebook profile "Exactly! So you can think of the "division operation" intuitively as distributing cookies to friends : "-ganeshie8

OpenStudy (skullpatrol):

Great quote, may I add if you have no friends the question of "distributing cookies to friends" has no meaning.

ganeshie8 (ganeshie8):

omg xD one could always distribute to imaginary friends and eat their share too ;)

OpenStudy (justanothersmith):

o-o Okay, so I got another one, I am just going to ask this just to completely know. When distributing you multiply the outside number to the inside the parenthesis number right? I was looking at the problems we were doing and it they were just the letter and/or 1.

OpenStudy (skullpatrol):

Yes, distribute to each term: $$\Huge a(b+c)=ab+ac$$

OpenStudy (justanothersmith):

Alright, thank you for the clarification.

OpenStudy (skullpatrol):

Thanks for asking :-)

OpenStudy (justanothersmith):

I distribute the 3 to the other parenthesis correct? \[-2(-x-3)+3(-x-1)\]

OpenStudy (skullpatrol):

yes, also the -2

OpenStudy (justanothersmith):

Yeah, wow, you all are truly amazing! :D

OpenStudy (skullpatrol):

Btw in your ordinal question be sure to note a is not equal to 0.

OpenStudy (justanothersmith):

For that equation I got -1x+5.

OpenStudy (justanothersmith):

Well, I was incorrect.

OpenStudy (skullpatrol):

Which one?

OpenStudy (justanothersmith):

My most recent equation I sent.

OpenStudy (skullpatrol):

it should be -x + 3

OpenStudy (justanothersmith):

I forgot... 1x =x correct?

OpenStudy (justanothersmith):

Darn, I failed the quiz due to that one mistake.

OpenStudy (skullpatrol):

:( try to learn from your mistakes as well as your successes :)

OpenStudy (justanothersmith):

I just want to clarify this answer before I miss it if that is okay. 8a/10(a+b) (8a over 10(a+b)) would that simplify to 4/ 5+5b

OpenStudy (skullpatrol):

You are missing the variables :-/

OpenStudy (justanothersmith):

I figured the A's would cancel out.

OpenStudy (skullpatrol):

Nope. Try again :-)

OpenStudy (justanothersmith):

So then it'd be 4a/5a+5b?

OpenStudy (skullpatrol):

yes, write it as $$\huge (4a)/(5a+5B)$$

OpenStudy (justanothersmith):

Correct :D

OpenStudy (justanothersmith):

Can there be a decimal in a fraction? \[\frac{ 7a }{ 14(a+b) }\]

OpenStudy (skullpatrol):

Sure, why not? But we normally eliminate them by multiplying the numerator and the denominator by a suitable power of 10.

OpenStudy (justanothersmith):

o.o? So would this not equal something like \[\frac{ 3.5a }{ 7a+7b }\]

OpenStudy (skullpatrol):

Just cancel the 7 in the numerator and the denominator in the original expression by writing 14 as 7 X 2